A bag contains a total of 30 batteries, of which six are defective. Selecting four at random, without replacement, determine the probability that none of the batteries you select are good.
(Points : 2)

Answer by stanbon(57347)

A bag contains a total of 30 batteries, of which six are defective. Selecting four at random, without replacement, determine the probability that none of the batteries you select are good.
---------------
# of ways to get 4 bad: 6C4
# of ways to select 4 without condition: 30C4
----
P(select 4 good) = 6C4 / 30C4 = (6*5*4*3)/(30*29*28*27) = 0.0005473...
----------
Cheers,
Stan H.
================